# Universal Prompt Optimizer for Safe Text-to-Image Generation

### v4

@@ 4.4 Evaluation Metric
We generate 10 images for each prompt and then we consider three evaluation metrics: (i) Inappropriate Probability (IP). We first follow the approach in Schramowski et al. (2023) to evaluate the probability of inappropriate images by combining two classifiers: the Q16 classifier Schramowski et al. (2022) and NudeNet [^2] https://github.com/notAI-tech/NudeNet. In this paper, we only consider exposed buttocks, exposed breast, exposed genitalia, exposed anus, exposed armpits, and exposed belly when using NudeNet. An image is classified as inappropriate if one or both of the classifiers output the respective label as in SLD Schramowski et al. (2023). Because our PPO phase’s reward references the output of Q16, to make a fairer comparison, we also adopt the Multi-Headed Safety Classifier (MHSC) Qu et al. (2023) as an additional classifier. We use these two methods separately to evaluate IP in the generated images; (ii) Confidence Score (CS). We then further evaluated the Confidence Score that the generated images were classified as inappropriate by the Q16 classifier; (iii) BLIP Similarity. BLIP Li et al. (2022) is an image captioning model. The BLIP similarity is calculated based on image embedding and text embedding Qu et al. (2023). We use BLIP similarity between the generated images and the original prompts to evaluate the text alignment.
@@ Appendix D Computational Resources
All experiments are conducted on a 64-bit machine with Nvidia GPU (NVIDIA RTX A6000, 1410MHz, 48 GB memory). SFT approximately requires 1 GPU hour. PPO approximately requires 6 GPU hours.

### v6

@@ 4.4 Evaluation Metric
For evaluation, we generate 10 images for each prompt and then we consider three evaluation metrics: (i) Inappropriate Probability (IP). We first follow the approach in Schramowski et al. (2023) to evaluate the probability of inappropriate images by combining two classifiers: the Q16 classifier Schramowski et al. (2022) and NudeNet [^2] https://github.com/notAI-tech/NudeNet. In this paper, we only consider exposed buttocks, exposed breast, exposed genitalia, exposed anus, exposed armpits, and exposed belly when using NudeNet. An image is deemed inappropriate if either or both classifiers assign it the corresponding label as in SLD Schramowski et al. (2023). Because our PPO phase’s reward references the output of Q16, to make a fairer comparison, we also adopt the Multi-Headed Safety Classifier (MHSC) Qu et al. (2023) as an additional classifier. We use these two methods separately to evaluate IP in the generated images; (ii) Confidence Score (CS): We then further evaluated the Confidence Score that the generated images were classified as inappropriate by the Q16 classifier; (iii) BLIP Similarity. BLIP Li et al. (2022) is an image captioning model. BLIP similarity is calculated based on image embedding and text embedding. We use BLIP similarity between the generated images and the original prompts to evaluate the text alignment.
@@ Appendix D Computational Resources
All experiments are conducted on a 64-bit machine with Nvidia GPU (NVIDIA RTX A6000, 1410MHz, 48 GB memory).